Please show your work.

3. Graph the following budget set. The two goods in question are food and clothing. Suppose that the individual has 500 dollars that can be spent on either food or clothing and that a 200 dollar subsidy is available for food and a 150 dollar subsidy is available for clothing. The price of a unit of food is 10 dollars and the price of a unit of clothing is 25 dollars. The intercepts, the slopes of the budget set, and any points where the slopes of the budget set changes must be labeled on the graph. This question is worth 10 points.

Please show your work.
Answer #1

According to the question,

Income of the consumer - 500$

Two commodities - Food and clothing

Subsidy:

Food - 200$

Clothing- 150$

Price of Food - 10$

Price of Clothing - 25$

Budget line = pf* F+ pc* C = Income

10*F + 25*C= 500

The slope of budget line = -pf/pc = -10/25= -.40

rW4Pu4TUOD04eEh4HCGIkCAAAECBIILOLcOTmpAA

What will happen to the budget line when a consumer receives subsidy:

tZaLHDZXa5oAAAAASUVORK5CYII=

There is a parallel shift f budget line form DE to AB with kink. A slope of the new budget line AB remains the same. However, the intercept changes for both the goods. No matter how he allocates his income but consumer have 150 $ to spent on clothing which can get him 6 units of cloth and 200$ for food from which he can get 20 units of food. It is a kinked demand curve where kink is at both ends.